# Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference

### v2

@@ A.3 Proof of Lemma 1
$\quad\quad\quad\quad\quad\cdot\mathbb{E}[\mathbb{E}[(Y_{i}-\mu(\bar{a}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta}_{0}))\mid\mathbf{X}^{out}_{i},\bar{A}_{i},\mathbf{H}_{i}]]]$
@@ A.3.1 Proof for Theorem 1
  1. (A11)
    As $n,J\rightarrow\infty$, $\frac{J}{n}\rightarrow\mathcal{R}$ a fixed constant.
  2. (A12)
    $\Sigma_{\beta,\alpha}$, $\Sigma_{\phi}$ and $\Sigma_{\gamma}$, defined below in Section A.3.1, is finite and has a non degenerate limit
  3. (A13)
    $\Omega_{\psi}$, defined below in Section A.3.1, is positive semi-definite.
  4. (A14)
    Regularity of parameter space. Parameters $\boldsymbol{\gamma}$ lie in a compact space $\boldsymbol{\gamma}$ with $\boldsymbol{\gamma}_{0}$ as an inner point.
  5. (A15)
    The score function for $A_{j}$, $\phi_{\gamma}$, is continuously differentiable in $\mathbf{X}^{int}_{j}$ and $\boldsymbol{\gamma}$, and dominated by an integrable function.
  6. (A16)
@@
Assumptions (A12)-(A13) are for a well-defined covariance matrix similar to (A9)-(A10) in Section 2.3.1. Assumptions (A14)-(A17) are primarily for showing Lemma 2, which shows we have asymptotic normality of estimating $\boldsymbol{\gamma}_{0}$. We still require (A7)-(A8) for consistency of our estimates.
@@
Let $\Phi_{n}^{\mathrm{aug}}$ be $\Phi_{n}$ augmented with Equation 11. For example, if $f_{0}$ is linear in $\mathbf{X}^{out}_{i}$, we will augment with $(1,\mathbf{X}^{out}_{i})$. I.e.
$\Phi_{n}^{\mathrm{aug}}$ $=\begin{bmatrix}\sum_{i=1}^{n}&\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))\\ \sum_{i=1}^{n}\sum_{j=1}^{J}&\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta})))\\ &\quad\quad\quad\quad\quad\cdot\frac{1}{J}H_{ij}(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))\end{bmatrix}$ (A.1)
@@
$\Phi_{n}^{\mathrm{aug}}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\gamma}})$ $=\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})+\frac{\partial\Phi_{n}^{\mathrm{aug}}(\boldsymbol{\beta},\boldsymbol{\alpha},\boldsymbol{\gamma}_{0})}{\partial(\boldsymbol{\beta}^{\top},\boldsymbol{\alpha}^{\top})}|_{\boldsymbol{\beta}=\boldsymbol{\beta}_{0},\boldsymbol{\alpha}=\boldsymbol{\alpha}_{0}}\begin{bmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\\ \hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\end{bmatrix}$
@@
Under regularity conditions (A7)-(A8) and (A14)-(A15), we have consistency of $\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\gamma}}$ and boundedness of these derivatives. Hence, $T_{2}=R_{nJ}=o_{P}(\sqrt{n})$, noting that remainder terms corresponding to $\boldsymbol{\gamma}$ are $o_{P}(\sqrt{J})=o_{P}(\sqrt{n})$ by (A11).
@@
$\phi_{i}^{\mathrm{aug}}$ $=\begin{bmatrix}&\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(A;\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(A;\mathbf{X}^{out}_{i},\mathbf{H}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))\\ &\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(A;\mathbf{X}^{out}_{i},\mathbf{H}_{i},\boldsymbol{\alpha},\boldsymbol{\beta})))\\ &\quad\quad\quad\quad\quad\cdot\frac{1}{J}\sum_{j=1}^{J}H_{ij}(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))\end{bmatrix}$ (A.2)
@@
Note that conditioning on $\mathbf{X}^{out}_{i},\mathbf{X}^{int}_{j},\mathbf{H}_{i},A$, each term on the RHS is mean 0.
@@
$\Omega_{\phi}=(\Sigma_{\beta,\alpha}^{-1})\Sigma_{\phi}(\Sigma_{\beta,\alpha}^{-1})^{\top}$ (A.3)
$\Omega_{\gamma}=(\frac{1}{\sqrt{\mathcal{R}}}\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma})\Omega_{\varepsilon}(\frac{1}{\sqrt{\mathcal{R}}}\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma})^{\top}$ (A.4)
@@
$\Omega=\Omega_{\phi}+\Omega_{\gamma}$ (A.5)
@@
A.4 Propensity Score Helper Lemma
@@
For Theorem 1, we require consistent estimation and asymptotic normality of propensity score model $\pi$. We state this in the following helper lemma.
@@
Assume (A7), (A11), (A14)-(A17). Then we have consistent estimation and asymptotic normality of propensity score model $\pi$. I.e.
[…]
@@
A.5 Additional Helper Results
The following is Lemma 10.5 from (Kosorok, 2008), restated here for convenience.
@@ A.5 Additional Helper Results
[…]
Appendix B Descriptive Statistics
@@
Table B.1: Summary of intervention level covariates from power plant data.
@@
Table B.2: Summary of outcome level covariates from Medicare Beneficiary data. The total number of IHD Hospitalizations in the dataset (2005) are 600,346 and the total rate of IHD Hospitalizations per 10k person-years is 775.362.
Appendix C Additional Simulation Details
C.1 Simulation Parameters
@@ C.2 Simulation Metrics
$\texttt{Bias}=||\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}_{0}||_{2}$ (C.1)
and the root mean squared error in the estimated treatment effect of interest is given by
$\texttt{RMSE}=\mbox{RMSE}(\widehat{\texttt{TotalEffect}},\texttt{TotalEffect})$ (C.2)
Appendix D Data Preprocessing
D.1 Intervention Unit Trimming
We trim intervention units with low treatment probabilities. The histogram for fitted intervention unit probabilities are shown in Section D.1.
D.2 Cost Model Training
We select our model based on Normalized Mean Absolute Error (NMAE). Table D.1 shows the NMAE for each cost model. For a model $\mathcal{M}$ with predictions $\hat{C}_{j}=\mathbb{E}[C_{j}\mid\mathbf{X}^{int}_{j};\mathcal{M}]$, NMAE is calculated as follows:
$NMAE(\mathcal{M})=\frac{1}{J}\sum_{j=1}^{J}\frac{|C_{j}-\hat{C}_{j}|}{|\hat{C}_{j}|}$ (D.1)
@@ D.2 Cost Model Training
Table D.1: Cost model NMAE on the validation set.
The random forest model performs best on our validation dataset, so random forest is used to generate cost predictions. The variable importance plot (under Increase in Node Purity) is shown in Section D.2. We find Total SO2 Emissions and and Total Load yield high variable importance for predicting cost, which makes sense given the scientific domain.
In Fig. D.1, we display the random forest model predictions against the true values, on the training and validation set.
@@
Figure D.1: (a) displays the predictions against truth in our training set and (b) displays the predictions against the truth in our validation set
Appendix E Additional Analysis Details
E.1 Exploratory Analysis of $\widehat{\texttt{TotalEffect}}$
We perform an exploratory analysis of $\texttt{TotalEffect}_{j}$ testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$. The p-values for each of the $J$ power plants are shown in Fig. E.1. Approximately 95% of p-values are $<0.05$. Caution should be taken in interpreting these findings, as dependence and false discovery should be accounted for in a formal statistical test. Regardless, these findings support that there are statistically significant population treatment effects with each scrubber installation.
Figure E.1: P-Values from testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$.
E.2 Nonlinear Policy Analysis
In Section 4, we performed policy analysis using linear models akin to Equation 15, Equation 16. In this section, we present a brief overview of results using nonlinear models. As shown in Table E.1, the potential reduction in hospitalization rates and counts increases notably with more complex models. As stated in the main text, we report the linear model results for interpretability and to err on the conservative side.
@@ E.2 Nonlinear Policy Analysis
Linear 44.51 31,906
Quadratic 58.50 58,984
Cubic 66.58 76,761
Sin/Cos 67.28 67,388
Table E.1: Estimated reductions in IHD hospitalization rates per 10k person-years and IHD hospitalization counts, varying the model link, with no budget constraints.
E.3 Count Analysis
The following are figures akin to Fig. 2, but for IHD Hospitalization Counts. The counts are obtained by multiplying the rates by the number of person-years observed for each zip code. As shown in Fig. E.2, $>99\%$ of power plants are protective. Many of the plants that are most protective in terms of rates are also protective in terms of counts. However, there are more plants that are similarly protective in the northeast and southeastern US. Similar to the cost-constrained rates analysis, much of the gain is experienced from 10% – 30% of universal scrubber treatment cost. In addition, from 60%, much of the gain stabilizes.
@@ E.3 Count Analysis
Figure E.2: (a) displays the plot of $\widehat{\texttt{TotalEffect}}_{j}$ on a US map, colored by intensity of $\widehat{\texttt{TotalEffect}}_{j}$. (b) displays the histogram of $\widehat{\texttt{TotalEffect}}_{j}$.

### v3

@@ 9 Supplementary Material
$\quad\quad\quad\quad\quad\cdot\mathbb{E}[\mathbb{E}[(Y_{i}-\mu(\bar{a}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\theta}_{0}))\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]]$
@@
  1. (R1’)
[…]
  2. (R2’)
    Regularity of parameter space. Parameters $(\boldsymbol{\gamma},\boldsymbol{\theta})$ lie in a compact space $\mathbf{\Gamma\cup\Theta}$ with $(\boldsymbol{\gamma}_{0},\boldsymbol{\theta}_{0})$ as an inner point.
  3. (R3’)
    $\Sigma_{\beta,\alpha}$, $\Sigma_{\phi}$ and $\Sigma_{\gamma}$, defined below in Section Appendix J.4, is finite and has a non-degenerate limit
  4. (R4’)
    $\Omega_{\psi}$, defined below in Section Appendix J.4, is positive semi-definite.
  5. (R5’)
@@
  7. (R7’)
    As $\rightarrow\infty, Jn\frac{J}{n}\rightarrow\mathcal{R}$ a fixed constant.
Assumptions (R3)-(R4) are for a well-defined covariance matrix similar to (R3)-(R4) in Section 2.3.2. Assumptions (R5)-(R6) in Section 2.3.2 are primarily for showing Lemma 2, which shows we have asymptotic normality of estimating $\boldsymbol{\gamma}_{0}$.
@@
Let $\Phi_{n}^{\mathrm{aug}}$ be $\Phi_{n}$ augmented with Eq. 10. For example, if $f_{0}$ is linear in $\mathbf{X}^{out}_{i}$, we will augment with $(1,\mathbf{X}^{out}_{i})$. I.e.
$\Phi_{n}^{\mathrm{aug}}$ $=\begin{bmatrix}\sum_{i=1}^{n}&\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))\\ \sum_{i=1}^{n}\sum_{j=1}^{J}&\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(\bar{A}_{i};\mathbf{X}^{out}_{i},\boldsymbol{\alpha},\boldsymbol{\beta})))\\ &\quad\quad\quad\quad\quad\cdot\frac{1}{J}H_{ij}(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))\end{bmatrix}$ (Appendix J.1)
@@
[…]
@@
Under regularity conditions (R1)-(R4), we have consistency of $\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\gamma}}$ and boundedness of these derivatives. Hence, $T_{2}=R_{nJ}=o_{P}(\sqrt{n})$, noting that remainder terms corresponding to $\boldsymbol{\gamma}$ are $o_{P}(\sqrt{J})=o_{P}(\sqrt{n})$ by (R7).
@@
$\phi_{i}^{\mathrm{aug}}$ $=\begin{bmatrix}&\frac{\partial}{\partial\boldsymbol{\alpha}}f_{0}(A;\mathbf{X}^{out}_{i},\boldsymbol{\alpha})(Y_{i}-\mu(A;\mathbf{X}^{out}_{i},\mathbf{H}_{i},\boldsymbol{\alpha},\boldsymbol{\beta}))\\ &\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})(Y_{i}-\mu(A;\mathbf{X}^{out}_{i},\mathbf{H}_{i},\boldsymbol{\alpha},\boldsymbol{\beta})))\\ &\quad\quad\quad\quad\quad\cdot\frac{1}{J}\sum_{j=1}^{J}H_{ij}(A_{j}-\pi(\mathbf{X}^{int}_{j},\boldsymbol{\gamma}))\end{bmatrix}$ (Appendix J.2)
@@
Note that conditioning on $\mathbf{X}^{out}_{i},\mathbf{X}^{int}_{j},\mathbf{H}_{i},\mathbf{A}$, each term on the RHS is mean 0.
@@
$\Omega_{\phi}=(\Sigma_{\beta,\alpha}^{-1})\Sigma_{\phi}(\Sigma_{\beta,\alpha}^{-1})^{\top}$ (Appendix J.3)
$\Omega_{\gamma}=(\frac{1}{\sqrt{\mathcal{R}}}\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma})\Omega_{\varepsilon}(\frac{1}{\sqrt{\mathcal{R}}}\Sigma_{\beta,\alpha}^{-1}\Sigma_{\gamma})^{\top}$ (Appendix J.4)
@@
$\Omega=\Omega_{\phi}+\Omega_{\gamma}$ (Appendix J.5)
@@
Appendix J.5 Proof for Theorem 4
Proof.
Below, proofs will be conditional on $\mathbf{H}_{i}$. By Jensen’s, it suffices to bound $|\mathbb{E}[\hat{\eta}_{j}-\eta_{j}]|$.
$\mathbb{E}\frac{1}{n}(\hat{\eta}_{j}-\eta_{j})$ $=\mathbb{E}_{n}[H_{ij}(f_{A}(\mathbf{X}^{out}_{i},\hat{\boldsymbol{\beta}})-f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}))]$
$=\mathbb{E}_{n}[H_{ij}(f_{A}(\mathbf{X}^{out}_{i},\hat{\boldsymbol{\beta}})-\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\hat{\boldsymbol{\beta}}))+H_{ij}(\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\hat{\boldsymbol{\beta}})-f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}))]$
$\leq\mathbb{E}\sup_{\boldsymbol{\beta}_{s}}\frac{1}{n}\sum_{i}(H_{ij}(f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})-\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})))$
where the 3rd line follows from our unbiased estimation equation. We choose to bound this quantity using symmetrization and spatial mixing assumptions combined with independent blocking from Bernstein [1927]. In this way, we handle the spatial dependence and present an argument easily extensible to more general, nonparametric function classes.
We formalize our spatial setup as in Jenish and Prucha [2009]. Let $D_{n}\subseteq D\subseteq\mathbb{R}^{2}$ for $D_{n}$ a finite subset. $\mathbf{X}^{out}_{i}$ lie in $D_{n}$. Define $L(i)$ as the location function; $L(i)$ returns the 2d coordinate location of unit $i$ in $D$. Take metric $\rho(i,i^{\prime})=\max_{1\leq k\leq 2}|L(i)_{k}-L(i^{\prime})_{k}|$, which induces a norm $|i|=\max_{1\leq k\leq 2}|L(i)_{k}|$, for $L(i)_{k}$ denoting the $k$-th component of the coordinate location of unit $i$. The distance between two subsets $U,V\subset D$ is defined as $\rho(U,V)=\inf\{\rho(i,i^{\prime}):i\in U,i^{\prime}\in V\}$, and the cardinality of a finite subset $U$ of $D$ is denoted $|U|$. Further define the spatial mixing coefficient (overloading notation on $\beta$):
For $U\subset D_{n},V\subset D_{n}$, define $\sigma_{n}(U)=\sigma(\mathbf{X}^{out}_{i}:i\in U)$, the sigma field over $U$, and $\beta_{n}(U,V)=\beta(\sigma_{n}(U),\sigma_{n}(V))=\frac{1}{2}\sup\sum_{U^{\prime}_{s}}\sum_{V^{\prime}_{r}}|\mathbb{P}(U^{\prime}_{s}\cap V^{\prime}_{r})-\mathbb{P}(U^{\prime}_{s})\mathbb{P}(V^{\prime}_{r})|$ where $\{U^{\prime}_{s}\},\{V^{\prime}_{r}\}$ are finite partitions of $U$ and $V$ respectively [Bradley, 2005]. Then $\beta$ mixing coefficient on $D_{n}$ is given by:
$\beta_{s,s^{\prime}}(r)=\sup\{\beta_{n}(U,V):|U|\leq s,|V|\leq s^{\prime},\rho(U,V)\geq r\}$
Also define $\overline{\beta}_{s,s^{\prime}}(r)=\sup_{n}\beta_{s,s^{\prime},n}(r)$ to account for sampling region variability.
Spatial Assumptions
  1. (S1)
    Infinite Sampling Regime: $D$ is infinitely countable and all elements are at least $\rho_{0}>1$ apart for all $i,i^{\prime}\in D$.
  2. (S2)
    Spatial Mixing: For $\tau>2$, $r\overline{\beta}_{s,s^{\prime}}(r)\leq(s+s^{\prime})\cdot r^{-\tau}$
These assumptions are standard in spatial statistics literature [Jenish and Prucha, 2009]. With the setup from (S1)-(S2)), we are equipped to bound this quantity.
First partition the grid $D_{n}$ into blocks with area $d_{n}^{2}$ yielding $L_{n}$ blocks. For $z\in\{0,1\}^{2}$, let $type(i)$ be whether or not sample $i$ lies in some block $l$ corresponding to ‘type z’, or spatial locations with block $l$ having the 1st coordinate being even/odd (0/1) and the second coordinate being even/odd (0/1). This permits us to employ the independent blocking from Bernstein on $\mathcal{B}_{z,l}=\{\mathbf{X}^{out}_{i}\in\mathcal{B}_{l}\mbox{ and }type(i)=z,\forall l\}$, or the data samples that are of type $z$ and lie in $\mathcal{B}_{l}$. We follow the strategy from Kuznetsov and Mohri [2017]: We first show that the average deviation can be bounded in terms of independent blocks, and thus i.i.d. tools can be readily applied. Mcdiarmid’s inequality and Rademacher averages along with Dudley’s integral will be employed to control block function complexities.
Bound on Independent Block Copy in terms of mixing quantities. Let $\mathcal{B}_{z,l}^{\prime}$ be independent copies of $\mathcal{B}_{z,l}$, and $\mathcal{B}_{z}$ be the blocks $\mathcal{B}_{z^{\prime},l}$ with $z=z^{\prime}$ (and $\mathcal{B}_{z}^{\prime}$ similarly). By Proposition 1 of Kuznetsov and Mohri [2017], we have for an $M$-bounded function $h$ and fixed $z$,
$\mathbb{E}[h(\mathcal{B}_{z}^{\prime})-h(\mathcal{B}_{z})]\leq L_{n}\cdot|D_{n}|\sup_{s,s^{\prime}}\overline{\beta}_{s,s^{\prime}}(d_{n})$
Take $h$ to be the block averages for block type $z$, or $h(\mathcal{B}_{z})=\mathbb{E}[\frac{1}{|\mathcal{B}_{z}|}\sum_{l}\sum_{i\in\mathcal{B}_{z,l}}H_{ij}(f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})-\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s}))]$. Then it follows from the union bound that
$\mathbb{P}[\frac{1}{n}\sum_{i=1}^{n}H_{ij}$ $(f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})-\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})>\delta]$
$\leq\sum_{z}\mathbb{P}[h(\mathcal{B}_{z}^{\prime})-\mathbb{E}h(\mathcal{B}_{z}^{\prime})>\delta-\mathbb{E}h(\mathcal{B}_{z}^{\prime})]+L_{n}\cdot\sup_{s,s^{\prime}}\overline{\beta}_{s,s^{\prime}}(d_{n})$
Now operating on these independent blocks, we can prove deviation inequalities using tools from the independently distributed setting. In particular, we apply a standard concentration bound and then an empirical process bound.
Concentration Inequality. First, a one block difference yields an empirical deviation of $M\frac{d_{n}}{n}$. By Mcdiarmid’s inequality,
$\mathbb{P}[\frac{1}{|\mathcal{B}_{z}^{\prime}|}\sum_{i\in\mathcal{B}_{z}^{\prime}}H_{ij}(f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})-\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})>\delta]$ $\leq\exp{(\frac{-2|D_{n}|^{2}(\delta+\mathbb{E}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s}))^{2}}{L_{n}d_{n}^{2}M^{2}})}$
Empirical Process Bound. It remains to bound $\mathbb{E}h(\mathcal{B}_{z}^{\prime})$. By symmetrization and Dudley’s integral,
$\mathbb{E}h(\mathcal{B}_{z}^{\prime})$ $\leq\mathbb{E}\sup_{\boldsymbol{\beta}_{s}}\frac{1}{|\mathcal{B}_{z}^{\prime}|}\sum_{l}\sum_{i\in B_{z,l}^{\prime}}\sigma_{i}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta}_{s})$
$\lesssim\frac{1}{\sqrt{n}}$
for $\sigma_{i}$ i.i.d. Rachemacher r.v.’s. The last line follows since $\boldsymbol{\beta}_{s}$ is finite-dimensional.
Integrating out the tail and employing (S2), we have :
$\mathbb{E}_{n}[\hat{\eta}_{j}-\eta_{j}]\lesssim\sqrt{\frac{1}{|D_{n}|}}+d_{n}\sqrt{\frac{1}{|D_{n}|}}$
We can choose $d_{n}$ such that $\frac{d_{n}}{|D_{n}|}=o_{p}(1)$, or $d_{n}$ grows slower than $D_{n}$. Noting that we took $\rho_{0}=1$, the result follows.
This (and our Section 2.3.2, Theorem 3) can be easily extended to nonparametric regression models to encompass splines, polynomial, fourier, and other bases. If we consider nonparametric models, we require our function class to be Donsker and formalized conditions on the approximation error from the optimal linear projection onto $k$ bases Belloni et al. [2015].
Appendix J.6 Proof for Theorem 5
Proof.
$\mathbb{E}[Y_{i}(\hat{\pi}_{j}^{*})-Y_{i}(\pi_{j}^{*})]$ $=\mathbb{E}[\frac{1}{J}\sum_{j=1}^{J}H_{ij}f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})(\hat{\pi}_{j}^{*}-\pi_{j}^{*})]$
$=\mathbb{E}[\frac{1}{J}\sum_{j=1}^{J}H_{ij}f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})[\textbf{1}\{\hat{\eta}_{j}<0\}-\textbf{1}\{\eta_{j}<0\}]$
$=\frac{1}{J}\sum_{j=1}^{J}H_{ij}\mathbb{E}[f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})\textbf{1}\{\hat{\eta}_{j}<0\}]-\mathbb{E}[f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})\textbf{1}\{\eta_{j}<0\}]$
We consider the inner expectation:
$\mathbb{E}[f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})\textbf{1}\{\hat{\eta}_{j}<0\}]$ $-\mathbb{E}[f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})\textbf{1}\{\eta_{j}<0\}]$
$=\mathbb{E}[f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})\mid\textbf{1}\{\hat{\eta}_{j}<0\}]\cdot\mathbb{P}[\hat{\eta}_{j}<0]-\mathbb{E}[f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta})\mid\eta_{j}<0]\cdot\mathbb{P}[\eta_{j}<0]]$
By Theorem 4, if $\eta_{j}<0$, $\textbf{1}\{\hat{\eta}_{j}<0\}$ occurs with high probability (converging at $\sqrt{n}$ rate). Thus, our estimated policy converges to the optimal policy.
Appendix J.7 Propensity Score Helper Lemma
@@
For Theorem 3, we require consistent estimation and asymptotic normality of propensity score model $e$. We state this in the following helper lemma.
@@
Assume the setup from Theorem 3. Then we have consistent estimation and asymptotic normality of propensity score model $\pi$. I.e.
$\sqrt{J}(\hat{\boldsymbol{\gamma}}-\boldsymbol{\gamma}_{0})\xrightarrow{d}N(0,\Omega_{\varepsilon})$
@@
Appendix J.8 Additional Helper Results
The following is Lemma 10.5 from [Kosorok, 2008], restated here for convenience.
@@
Let $Z_{1}\dots Z_{n}$ be i.i.d. euclidean random vectors with $\mathbb{E}Z=0$ and finite second moment, independent of i.i.d. sequence of real random variables $\zeta_{1}\dots\zeta_{n}$ mean 0 and variance 1. Then, conditionally on $Z_{1}\dots Z_{n}$, $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\zeta_{i}Z_{i}\xrightarrow{d}N(0,\mathrm{Cov}(Z))$ for almost all sequences $Z_{1}\dots Z_{n}$
Appendix K Descriptive Statistics
@@
Table Appendix K.1: Summary of intervention level covariates from power plant data.
@@
Table Appendix K.2: Summary of outcome level covariates from Medicare Beneficiary data.
Appendix L Additional Simulation Details
Appendix L.1 Simulation Parameters
@@
[…]
and the RMSE
$\texttt{RMSE}=\mbox{RMSE}(\widehat{\texttt{TotalEffect}},\texttt{TotalEffect})$ (Appendix L.2)
Appendix M Data Preprocessing
Appendix M.1 Intervention Unit Trimming
We trim intervention units with low treatment probabilities. The histogram for fitted intervention unit probabilities are shown in Fig. Appendix M.1.
Figure Appendix M.1: Plot of the estimated intervention unit propensities.
Appendix M.2 Cost Model Training
We select our model based on Normalized Mean Absolute Error (NMAE). For a model $\mathcal{M}$ with predictions $\hat{C}_{j}=\mathbb{E}[C_{j}\mid\mathbf{X}^{int}_{j};\mathcal{M}]$, the NMAE is calculated as follows.
$NMAE(\mathcal{M})=\frac{1}{J}\sum_{j=1}^{J}\frac{|C_{j}-\hat{C}_{j}|}{|\hat{C}_{j}|}$ (Appendix M.1)
Table Appendix M.1 shows the NMAE for each cost model.
@@
Table Appendix M.1: Cost model NMAE on the validation set.
The random forest model performs best on our validation dataset, so random forest is used to generate cost predictions. The variable importance plot (under Increase in Node Purity) is shown in Fig. Appendix M.2. We find Total SO2 Emissions and and Total Load yield high variable importance for predicting cost, which aligns with the scientific domain.
Figure Appendix M.2: Plot of the variable importances for our cost model.
In Fig. Appendix M.3, we display the random forest model predictions against the true values, on the training and validation set.
@@
Figure Appendix M.3: (a) displays the predictions against truth in our training set and (b) displays the predictions against the truth in our validation set.
Appendix N Additional Analysis Details
Appendix N.1 Exploratory Analysis of $\widehat{\texttt{TotalEffect}}$
We perform an exploratory analysis of $\texttt{TotalEffect}_{j}$ testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$. The p-values for each of the $J$ power plants are shown in Fig. Appendix N.1. Approximately 70% of p-values are <0.05<0.05. Caution should be taken in interpreting these findings, as dependence and false discovery should be accounted for in a formal statistical test. Regardless, these findings support that there are statistically significant population treatment effects with each scrubber installation.
Figure Appendix N.1: P-Values from testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0$.
Appendix N.2 Nonlinear Policy Analysis
In Section 4, we performed policy analysis using linear models akin to Section 3. In this section, we present a brief overview of results using nonlinear models. As shown in Table Appendix N.1, the potential reduction in hospitalization rates and counts do not increase with the complexity.
@@
Linear 55.30 38,372
Quadratic 33.58 22,309
Cubic 45.87 31,958
Sin/Cos 67.28 50,990
Table Appendix N.1: Estimated reductions in IHD hospitalization rates per 10k person-years and IHD hospitalization counts, varying the model link, with no budget constraints.
Appendix N.3 Count Analysis
The following are figures akin to Fig. 2, but for IHD Hospitalization Counts. The counts are obtained by multiplying the rates by the number of person-years observed for each zip code. There are more plants that are similarly protective in the northeast and southeastern US. Similar to the cost-constrained rates analysis, large gains are experienced from 10% – 30% of universal scrubber treatment cost. In addition, from 60%, much of the gain stabilizes.
@@
Figure Appendix N.2: Count Results. (a) displays the plot of $\widehat{\texttt{TotalEffect}}_{j}$ on a US map, colored by intensity of $\widehat{\texttt{TotalEffect}}_{j}$. (b) displays the histogram of $\widehat{\texttt{TotalEffect}}_{j}$.